# Optical design and sensitivity optimization of Cryogenic sub-Hz cROss torsion bar detector with quantum NOn-demolition Speed meter (CHRONOS)

### v1

@@ Abstract
We present the optical design and sensitivity modeling of the 2.5 m Cryogenic sub-Hz cROss torsion-bar detector with quantum NOn-demolition Speed meter (CHRONOS), a triangular Sagnac speed-meter interferometer incorporating power- and signal-recycling techniques. Using ABCD-matrix analysis and Finesse3 simulations, we show that stable eigenmodes are obtained with optimized mirror curvatures and focal placements, achieving mode-matching efficiencies above 99.5 The resulting configuration reaches a quantum-noise-limited strain sensitivity of $h\simeq 3\times 10^{-18},\mathrm{Hz^{-1/2}}$ at 1 Hz, with a ring-cavity finesse $\mathcal{F}\simeq 3.1\times 10^{4}$ and round-trip Gouy phase $\psi_{\mathrm{rt}}\approx 153^{\circ}$. The power-recycling cavity detuning ($\phi_{p}=-85^{\circ}$) dominates the low-frequency quantum noise, while the signal-recycling cavity detuning ($\phi_{s}=0^{\circ}$) mainly introduces a uniform quadrature rotation. The optimal homodyne angle ($\zeta_{\mathrm{opt}}\simeq 46^{\circ}$) balances shot-noise and radiation-pressure effects to give the best sensitivity near 1 Hz. Assuming end-mirror reflectivity $R_{\mathrm{ETM}}=99.9999\%$ under cryogenic operation at 10 K, CHRONOS can achieve quantum-noise-limited performance on a laboratory scale and serve as a testbed for future long-baseline, cryogenic interferometers probing sub-hertz gravitational waves.
@@ I Introduction
Gravitational-wave astronomy has advanced rapidly through the success of ground-based interferometers such as LIGO, Virgo, and KAGRA Abbott et al. (2016, 2021a); Akutsu et al. (2019), and is being further extended into the millihertz band with the upcoming space-based Laser Interferometer Space Antenna (LISA) Amaro-Seoane et al. (2017a). However, a sensitivity gap remains in the sub-hertz band. Ground-based detectors are limited at low frequencies by seismic and environmental noise, while space missions cannot adequately cover this band due to constraints on arm length. Several terrestrial efforts specifically target this regime, such as the torsion-bar antenna (TOBA) Ando et al. (2010) and Torsion Pendulum Dual Oscillator (TorPeDO) McManus et al. (2017), both of which have demonstrated proof-of-principle operation of sub-hertz detectors.
@@
Opening the sub-hertz band would unlock new opportunities for gravitational-wave astronomy. In particular, the merger signals of intermediate-mass black hole (IMBH) binaries exhibit characteristic frequency evolution in this regime. For the Cryogenic sub-Hz cROss torsion bar detector with quantum NOn-demolition Speed meter (CHRONOS) sensitivity scale Inoue et al. (2025), a facility with an arm length of about 300 m could detect mergers with total mass $M\sim 10^{3}M_{\odot}$ out to 380 Mpc, enabling multi-band observations that bridge LISA and future ground-based detectors. Such observations would provide direct evidence for IMBHs and valuable insights into the formation pathways of supermassive black holes Amaro-Seoane et al. (2017b). Moreover, recent O4 observations have begun to reveal the black-hole mass distribution, with several candidate IMBH events already reported Abbott et al. (2021a, b). Measurements in the high-mass regime are particularly important for understanding black-hole formation scenarios, and CHRONOS will offer unique access to this parameter space.
@@
In addition, CHRONOS provides unique sensitivity to the stochastic gravitational-wave background ($\Omega_{\rm GW}$) of cosmological origin. With five years of integrated observation, the detector is expected to reach a level of $\Omega_{\rm GW}\sim 3\times 10^{-9}$ at 0.2 Hz. This would directly probe early-universe phenomena such as first-order phase transitions and cosmic strings, thus offering rare opportunities to test physics beyond the standard cosmological model Caprini et al. (2016); Ng et al. (2021).
@@
The broader scientific prospects and overall detector concept are discussed in detail in the companion CHRONOS design paper. In this study, we report on the first stage of development: the optical design and sensitivity evaluation of a 2.5 m-scale interferometer. The CHRONOS adopts a triangular Sagnac configuration with output recycling to enhance sensitivity. We investigate suitable sideband signals for controlling multiple cavities, and present design parameters for realistic optical configurations such as dual-recycling cavities. To achieve stable locking of the Sagnac cavity, we compute the mirror-curvature conditions required to reach the desired mode-matching efficiency (99.5%), and confirm consistency across the full optical design. Finally, we calculate sensitivity curves based on the concrete optical layout and discuss the fundamental sensitivity limits of the 2.5 m test facility.
@@ II Principle
The readout in this study follows the TOBA framework Ando et al. (2010), which is based on the torsional degree of freedom around the $z$ axis. The angular displacement $\theta$ of a bar-shaped test mass with moment of inertia $I$, torsional spring constant $\kappa$, and loss coefficient $\gamma$ obeys
@@
where the torque from the gravitational-wave tidal acceleration $a_{i}=\tfrac{1}{2}\,\ddot{h}_{ij}\,x^{j}$ is given by Saulson (1994); Forward (1978)
@@
$Q_{ij}\equiv\int_{V}\rho\!\left(x_{i}x_{j}-\tfrac{1}{3}r^{2}\delta_{ij}\right)dV.$ (3)
In the frequency domain, this becomes
$\theta(\Omega)$ $=\chi_{\theta}(\Omega)\!\left[-\frac{I_{\mathrm{eff}}}{2}\Omega^{2}\bigl(h_{+}F_{+}+h_{\times}F_{\times}\bigr)\right],$
@@
where the torsional resonance $\Omega_{t}=\sqrt{\kappa/I}$ is taken to be well below the observational band (sub-mHz to mHz). In the regime $\Omega\gg\Omega_{t}$, the susceptibility reduces to $\chi_{\theta}\simeq(-I\Omega^{2})^{-1}$, yielding
$\theta(\Omega)\simeq\frac{I_{\mathrm{eff}}}{2I}\bigl(h_{+}F_{+}+h_{\times}F_{\times}\bigr).$ (5)
The total and effective moments of inertia are
@@
such that $\eta\equiv I_{\mathrm{eff}}/I\simeq 1$ for an ideal thin rod. In the long-wavelength limit with $z$-axis incidence, the antenna patterns are Forward (1978)
@@
By combining two orthogonal bars (X and Y) into a differential readout $\theta=\theta_{X}-\theta_{Y}$, common-mode noise sources such as translational motion and laser intensity fluctuations are suppressed, while retaining sensitivity to linear combinations of the two polarizations Ando et al. (2010).
@@ III Optical layout
An overview of the optical layout is shown in Fig. 2. The experimental apparatus is based on a cross-torsion-bar configuration, with identical optical systems placed in the X and Y arms. After pre-stabilization, the input laser beam first passes through a pre-stabilized laser (PSL) system and an input mode cleaner (IMC), which stabilize and shape the spatial and frequency profiles Aasi et al. (2015); Abbott et al. (2021c). The beam is then split at the input beam splitter (IBS) and directed into the Sagnac interferometers of the X and Y arms. Within each arm, the light propagates through a triangular Sagnac ring cavity, and the returning beams are recombined at the output beam splitter (OBS). The phase difference between the two arms is extracted as the interferometric signal. For convenience, the following description focuses on the Y arm, but it should be noted that the X arm has an identical optical configuration.
In the overall CHRONOS program, facility scales of 2.5 m, 40 m, and 400 m have been considered, and their sensitivities compared. In this paper, however, we focus on the 2.5 m-scale interferometer as the first step and discuss in detail the sensitivity evaluation Inoue et al. (2025).
@@
The central part of the interferometer is the triangular Sagnac ring cavity formed by the input test mass (ITM), the middle test mass (MTM), and the end test mass (ETM) located at the ends of each arm. In a simplified description, the closed ring path is defined by six mirrors including the ITM. To ensure proper mode matching, the ETMs and curved mirrors (CMs) are given finite radii of curvature, thereby defining an internal beam waist. The laser injection point is placed about 0.5 m away from the geometric center of the optics, and this asymmetric injection fixes the propagation path of the ring cavity. The Sagnac topology measures the velocity rather than the displacement of the test masses, thereby naturally suppressing radiation-pressure noise at low frequencies Chen (2003); Danilishin and Khalili (2012).
@@
Power-recycling mirror (PRM) is realized by a cavity formed with PRM1–PRM4 and the beam splitter (BS). This configuration enhances the circulating power inside the Sagnac ring without requiring excessively high input laser power. In particular, PR2 provides the main degree of freedom for mode matching: with a finite radius of curvature, it adjusts the external beam waist position and matches it to the internal eigenmode.
@@
At the output port, a signal-recycling mirror (SRM) is installed to control the extraction of gravitational-wave signal sidebands. By tuning the detuning of this cavity, the overall sensitivity curve—especially in the sub-hertz regime—can be reshaped Mizuno et al. (1993). Signal recycling also provides flexibility in optimizing quantum-noise suppression at low frequencies and sensitivity enhancement at high frequencies Kimble et al. (2001); Danilishin and Khalili (2012).
@@
An output mode cleaner (OMC) is placed at the detection port to remove higher-order spatial modes and unwanted carrier leakage before the light reaches the photodetectors to be -40 dB Aasi et al. (2015). This ensures that the detected field faithfully reflects the gravitational-wave sidebands of the interferometer. The final output is read out using balanced-homodyne detection (BHD). By interfering the interferometer output with a local oscillator on two photodiodes and subtracting the resulting photocurrents, arbitrary quadratures of the signal can be measured. Compared with conventional DC readout, this method improves quantum-noise performance at low frequencies Fricke et al. (2012); Vahlbruch et al. (2010).
@@ IV Cavity control
In this work, we adopt a triangular Sagnac topology that measures the velocity rather than the displacement of the test masses. This feature naturally suppresses radiation-pressure noise in the low-frequency band and is essential for extending the sensitivity toward the sub-hertz regime. The 2.5 m interferometer provides a realistic environment for experimentally validating optical stability, mode matching, control schemes, and quantum-noise readout strategies.
@@ IV.1 Sideband frequencies
These values reproduce the resonant and anti-resonant operating conditions originally established in the LIGO series Buonanno and Chen (2003), which define the optimal phase relationships between the power- and signal-recycling cavities to enhance interferometer response at specific frequency bands. While those configurations were first implemented in kilometer-scale detectors, the present design achieves equivalent optical conditions at the 2.5 m scale, demonstrating that the essential interferometric dynamics can be faithfully reproduced within a compact experimental platform.
Based on this configuration, three radio frequency (RF) sidebands are selected. Two of them are used for the main cavity control:
@@
This configuration provides sufficient discrimination for laser frequency stabilization while maintaining consistency with the overall interferometer frequency structure Kwee et al. (2012); Evans et al. (2015).
@@ IV.2 Control ports and control scheme
The CHRONOS interferometer employs a dual-recycled Sagnac configuration, using multiple detection ports and RF sidebands to stably control the common and differential degrees of freedom of the coupled cavities. This section describes the overall control scheme—including the initial alignment (INI) phase—along with the definition and optical roles of each control port Evans et al. (2015); Akutsu et al. (2021).
Initial alignment and INI control.
@@
The IMC is a four-mirror bow-tie cavity that serves both as a spatial mode filter and as a frequency reference Kwee et al. (2012); Akutsu et al. (2021). Each mirror is suspended by a single-pendulum suspension system, allowing isolation from ground vibrations and precise angular control. The cavity length is maintained by Pound-Drever-Hall (PDH) control Drever et al. (1983); Black (2001) using modulation at $f_{\mathrm{IMC}}=44.235\penalty 10000\ \mathrm{MHz}$. The reflected light is demodulated at the same frequency to produce an error signal, which drives coil-magnet actuators attached to the end mirror to control the length. This configuration follows the KAGRA-style IMC design Akutsu et al. (2021), which employs magnetic actuation of the suspended mirrors rather than piezoelectric (PZT) actuation, enabling large-stroke, low-frequency control with excellent stability. The transmitted beam serves as a monitor for the spatial mode and intensity stability, and its low-frequency component is also used for offset compensation in the laser-frequency control loop Kwee et al. (2012).
The RF phase reference of the IMC is phase-locked to the modulation sources at $f_{1}=29.5\penalty 10000\ \mathrm{MHz}$ and $f_{2}=58.980\penalty 10000\ \mathrm{MHz}$, ensuring a common demodulation phase for the Reflection port (REFL) and Pick-off port (POP) detection channels. Once this initial alignment phase is completed, the input beam with matched spatial mode, frequency, and phase is established, allowing transition to the green pre-lock and PRC/SRC control stages.
Reflection port (REFL).
The REFL port detects the light reflected from the input mirror of the PRC. This port directly monitors the total reflectivity of the interferometer and provides the primary error signal for the PRC length (PRCL). Demodulation of the reflected light at $f_{1}=29.5\penalty 10000\ \mathrm{MHz}$ produces the PRC error signal via the PDH technique Drever et al. (1983); Black (2001). In addition, the REFL port serves as an auxiliary monitor for the common Sagnac mode (CSAG), corresponding to the average optical path length of the clockwise (CW) and counter-clockwise (CCW) beams, and contributes to maintaining the overall optical resonance and laser-frequency stability.
Pick-off port (POP).
The POP port is a pick-off of the intracavity field in the PRC, designed to sense both the common and differential components of the optical field Aston et al. (2012). Two photodiodes are placed at positions with distinct Gouy phases to spatially separate optical modes:

### v2

@@ Abstract
We present the optical design and sensitivity modeling of the 2.5 m Cryogenic sub-Hz cROss torsion-bar detector with quantum NOn-demolition Speed meter (CHRONOS), a triangular Sagnac speed-meter interferometer incorporating power- and signal-recycling techniques. Using ABCD-matrix analysis and Finesse3 simulations, we show that stable eigenmodes are obtained with optimized mirror curvatures and focal placements, achieving mode-matching efficiencies above 99.5%. The configuration reaches a quantum-noise-limited strain sensitivity of $h\simeq 3\times 10^{-18}\,\mathrm{Hz^{-1/2}}$ at 1 Hz, with a ring-cavity finesse $\mathcal{F}\simeq 3.1\times 10^{4}$ and a round-trip Gouy phase $\psi\approx 153^{\circ}$. The power-recycling cavity detuning ($\phi_{p}=-85^{\circ}$) dominates the low-frequency quantum noise, while the signal-recycling cavity detuning ($\phi_{s}=0^{\circ}$) mainly introduces a uniform quadrature rotation. The optimal homodyne angle ($\zeta\simeq 46^{\circ}$) balances shot-noise and radiation-pressure effects to give the best sensitivity near 1 Hz. Assuming an end-mirror reflectivity $R_{\mathrm{ETM}}=99.9999\%$ under cryogenic operation at 10 K, CHRONOS can achieve quantum-noise-limited performance on a laboratory scale and serve as a testbed for future long-baseline cryogenic interferometers probing sub-hertz gravitational waves.
@@ I Introduction
Gravitational-wave astronomy has advanced rapidly through the success of ground-based interferometers such as LIGO, Virgo, and KAGRA Abbott2016 ; GWTC3 ; KAGRA2020 , and is being further extended into the millihertz band with the upcoming space-based Laser Interferometer Space Antenna (LISA) LISA2017 . However, a sensitivity gap remains in the sub-hertz band. Ground-based detectors are limited at low frequencies by seismic and environmental noise, while space missions cannot adequately cover this band due to constraints on arm length. Several terrestrial efforts specifically target this regime, such as the torsion-bar antenna (TOBA) Ando2010 and the Torsion Pendulum Dual Oscillator (TorPeDO) McManus2017TorPeDO , both of which have demonstrated proof-of-principle operation of sub-hertz detectors.
@@
Opening the sub-hertz band would unlock new opportunities for gravitational-wave astronomy. In particular, the merger signals of intermediate-mass black hole (IMBH) binaries exhibit characteristic frequency evolution in this regime. For the Cryogenic sub-Hz cROss torsion bar detector with quantum NOn-demolition Speed meter (CHRONOS) sensitivity scale inoue2025chronos , a facility with an arm length of about 300 m could detect mergers with total mass $M\sim 10^{3}M_{\odot}$ out to 340 Mpc, enabling multi-band observations that bridge LISA and future ground-based detectors. Such observations would provide direct evidence for IMBHs and valuable insights into the formation pathways of supermassive black holes AmaroSeoane2017 . Moreover, recent O4 observations have begun to reveal the black-hole mass distribution, with several candidate IMBH events already reported GWTC3 ; GWTC4pop . Measurements in the high-mass regime are particularly important for understanding black-hole formation scenarios, and CHRONOS will offer unique access to this parameter space.
@@
In addition, CHRONOS provides unique sensitivity to the stochastic gravitational-wave background (SGWB) $\Omega_{\rm GW}$ of cosmological origin. With five years of integrated observation, the detector is expected to reach a level of $\Omega_{\rm GW}\sim 3\times 10^{-9}$ at 0.2 Hz. This would directly probe early-universe phenomena such as first-order phase transitions and cosmic strings, thus offering rare opportunities to test physics beyond the standard cosmological model Caprini2016 ; Ng2021 ; inoue2025chronos .
@@
The broader scientific prospects and overall detector concept are discussed in detail in the companion CHRONOS design paper. In this study, we report on the first stage of development: the optical design and sensitivity evaluation of a 2.5 m-scale interferometer. The CHRONOS adopts a triangular Sagnac configuration with output recycling to enhance sensitivity. We investigate suitable sideband signals for controlling multiple cavities and present design parameters for realistic optical configurations such as dual-recycling cavities. To achieve stable locking of the Sagnac cavity, we compute the mirror-curvature conditions required to reach the desired mode-matching efficiency (99.5%), and confirm consistency across the full optical design. Finally, we calculate sensitivity curves based on the concrete optical layout and discuss the fundamental sensitivity limits of the 2.5 m test facility.
@@ II Principle
The readout in this study follows the TOBA framework Ando2010 , which is based on the torsional degree of freedom around the $z$ axis. The angular displacement $\theta$ of a bar-shaped test mass with moment of inertia $I$, torsional spring constant $\kappa$, and loss coefficient $\gamma$ obeys
@@
where the torque from the gravitational-wave tidal acceleration $a_{i}=\tfrac{1}{2}\,\ddot{h}_{ij}\,x^{j}$ is given by Saulson1994 ; Forward1978
@@
$Q_{ij}\equiv\int_{V}\rho\!\left(x_{i}x_{j}-\tfrac{1}{3}r^{2}\delta_{ij}\right)\,dV,$ (3)
where the dimension of the torsion bar is described in Fig. 1. In the frequency domain, this becomes
$\theta(\Omega)$ $=\chi_{\theta}(\Omega)\!\left[-\frac{I_{\mathrm{eff}}}{2}\Omega^{2}\bigl(h_{+}F_{+}+h_{\times}F_{\times}\bigr)\right],$ (4)
@@
where the torsional resonance $\Omega_{t}=\sqrt{\kappa/I}$ is taken to be well below the observational band (0.1–10 Hz). In the regime $\Omega\gg\Omega_{t}$, the susceptibility reduces to $\chi_{\theta}\simeq(-I\Omega^{2})^{-1}$, yielding
$\theta(\Omega)_{i}\simeq\frac{\eta_{g}}{2}\bigl(h_{+}F_{+}+h_{\times}F_{\times}\bigr),$ (6)
where $i=\{X,Y\}$. The total and effective moments of inertia are
@@
such that $\eta_{g}\equiv I_{\mathrm{eff}}/I=0.936$ for an ideal thin rod. In the long-wavelength limit with $z$-axis incidence, the antenna patterns are Forward1978
@@
By combining two orthogonal bars (X and Y) into a differential readout,
$\theta(\Omega)=\theta_{X}-\theta_{Y}=\eta_{g}\bigl(h_{+}F_{+}+h_{\times}F_{\times}\bigr),$ (9)
common-mode noise sources such as translational motion and laser intensity fluctuations are suppressed, while retaining sensitivity to linear combinations of the two polarizations Ando2010 .
@@ III Optical layout
An overview of the optical layout is shown in Fig. 2. The experimental apparatus is based on a cross-torsion-bar configuration, with identical optical systems placed in the X and Y arms. After pre-stabilization, the input laser beam first passes through a pre-stabilized laser system and an input mode cleaner (IMC), which stabilize and shape the spatial and frequency profiles aasi2015advanced ; abbott2021ligo . The beam is then split at the input beam splitter (IBS) and directed into the Sagnac interferometers of the X and Y arms. Within each arm, the light propagates through a triangular Sagnac ring cavity, and the returning beams are recombined at the output beam splitter (OBS). The phase difference between the two arms is extracted as the interferometric signal. For convenience, the following description focuses on the Y arm, but it should be noted that the X arm has an identical optical configuration.
In the overall CHRONOS program, interferometer facility scales of $2.5\penalty 10000\ \mathrm{m}$, $40\penalty 10000\ \mathrm{m}$, and $300\penalty 10000\ \mathrm{m}$ have been considered, and their sensitivities systematically compared in Ref. inoue2025chronos . These characteristic arm lengths were chosen to reflect existing and realistic test facilities, such as the ASGRAF prototype, the Glasgow $10\penalty 10000\ \mathrm{m}$ interferometer, the Caltech $40\penalty 10000\ \mathrm{m}$ interferometer, TAMA, and GEO 600, which span a representative range of laboratory- to observatory-scale systems. Although the arm length is varied across these configurations, the size and material properties of the test masses are assumed to be identical. In this paper, we therefore focus specifically on the $2.5\penalty 10000\ \mathrm{m}$-scale interferometer, as the first experimental step of CHRONOS, and discuss its optical design and sensitivity evaluation in detail.
The central element of the interferometer is a triangular Sagnac ring cavity formed by the input test mass (ITM), the middle test mass (MTM), and the end test mass (ETM) located at the ends of each arm. Two such ring cavities are installed symmetrically on the left and right sides of the torsion bar, and the angular rotation of the torsion bar is read out through the differential response of these cavities. In a simplified description, the closed Sagnac ring path is defined by six mirrors, including the ITMs.
@@
To ensure proper mode matching and stable cavity eigenmodes, the ETMs and curved mirrors (CMs) are assigned finite radii of curvature, thereby defining an internal beam waist within each ring cavity. The laser beam, split by the beam splitter (BS), is injected into the ring cavities via the power-recycling mirror (PRM); this asymmetric injection uniquely determines the propagation direction and optical path of the Sagnac rings. Owing to the Sagnac topology, the interferometer measures the velocity of the test masses, rather than their displacement, which naturally suppresses radiation-pressure noise at low frequencies chen2003 ; danilishin2012 .
@@
The PRM is implemented as a cavity formed by PRM1–PRM4 together with the BS. This configuration enhances the circulating optical power inside the Sagnac ring cavities without requiring excessively high input laser power. Among the power-recycling mirrors, PR2 plays a particularly important role in mode matching: by introducing a finite radius of curvature, PR2 adjusts the external beam waist position and matches it to the internal cavity eigenmode.
@@
At the output port, a signal-recycling mirror (SRM) is installed to control the extraction of gravitational-wave signal sidebands. By tuning the detuning of this cavity, the overall sensitivity curve—especially in the sub-hertz regime—can be reshaped mizuno1993 . Signal recycling also provides flexibility in optimizing quantum-noise suppression at low frequencies and sensitivity enhancement at high frequencies kimble2001 ; danilishin2012 .
@@
An output mode cleaner (OMC) is placed at the detection port to remove higher-order spatial modes and unwanted control sidebands before the light reaches the photodetectors, providing a suppression level of approximately $40\penalty 10000\ \mathrm{dB}$ aasi2015advanced . This filtering ensures that the detected optical field faithfully reflects the gravitational-wave signal sidebands of the interferometer.
The final output is read out using balanced-homodyne detection (BHD). In this scheme, the interferometer output is interfered with a local oscillator on two photodiodes, and the resulting photocurrents are subtracted, allowing an arbitrary optical quadrature of the signal to be measured. Compared with conventional DC readout, balanced-homodyne detection improves the quantum-noise performance, particularly at low frequencies fricke2012 ; vahlbruch2010 .
@@ IV Cavity control
In this work, we adopt a triangular Sagnac topology that measures the angular momentum rather than the angle of the test masses. This feature naturally suppresses radiation-pressure noise in the low-frequency band and is essential for extending the sensitivity toward the sub-hertz regime. The study of the 2.5 m interferometer provides a realistic environment for experimentally validating optical stability, mode matching, control schemes, and quantum-noise readout strategies.
Each reference port and sideband frequency is summarized in Table 1.
@@ IV.1 Sideband frequencies
These values reproduce the resonant and anti-resonant operating conditions originally established in LIGO, which define the optimal phase relationships between the power- and signal-recycling cavities to enhance interferometer response at specific frequency bands. While those configurations were first implemented in kilometer-scale detectors, the present design achieves equivalent optical conditions at the 2.5 m scale, demonstrating that the essential interferometric dynamics can be faithfully reproduced within a compact experimental platform.
Based on this configuration, three radio-frequency (RF) sidebands are selected. Two of them are used for the main cavity control:
@@
This configuration provides sufficient discrimination for laser frequency stabilization while maintaining consistency with the overall interferometer frequency structure Kwee2012 ; Evans2015 .
@@ IV.2 Control ports and control scheme
The CHRONOS interferometer employs a dual-recycled Sagnac configuration, using multiple detection ports and RF sidebands to stably control the common and differential degrees of freedom of the coupled cavities. This section describes the overall control scheme—including the initial alignment (INI) phase—along with the definition and optical roles of each control port Evans2015 ; akutsu2021kagra .
IV.2.1 Initial alignment and INI control
@@
The IMC is a four-mirror bow-tie cavity that serves both as a spatial mode filter and as a frequency reference Kwee2012 ; akutsu2021kagra . Each mirror is suspended by a single-pendulum suspension system, allowing isolation from ground vibrations and precise angular control. The cavity length is maintained by Pound–Drever–Hall (PDH) control Drever1983 ; Black2001 using modulation at $f_{\mathrm{IMC}}=44.235\penalty 10000\ \mathrm{MHz}$. The reflected light is demodulated at the same frequency to produce an error signal, which drives coil-magnet actuators attached to the end mirror to control the length. This configuration follows the KAGRA-style IMC design akutsu2021kagra , which employs magnetic actuation of the suspended mirrors rather than piezoelectric actuation, enabling large-stroke, low-frequency control with excellent stability. The transmitted beam serves as a monitor for the spatial mode and intensity stability, and its low-frequency component is also used for offset compensation in the laser-frequency control loop Kwee2012 .
The RF phase reference of the IMC is phase-locked to the modulation sources at $f_{1}=29.5\penalty 10000\ \mathrm{MHz}$ and $f_{2}=58.980\penalty 10000\ \mathrm{MHz}$, ensuring a common demodulation phase for the reflection port (REFL) and pick-off port (POP) detection channels. Once this initial-alignment phase is completed, an input beam with matched spatial mode, frequency, and phase is established, allowing a smooth transition to the green pre-lock stage and subsequent PRC/SRC control. The locations of the individual control readout ports are illustrated in Fig. 2.
IV.2.2 Reflection port (REFL)
The REFL port detects the light reflected from the input mirror of the PRC. This port directly monitors the total reflectivity of the interferometer and provides the primary error signal for the PRC length (PRCL). Demodulation of the reflected light at $f_{1}=29.5\penalty 10000\ \mathrm{MHz}$ produces the PRC error signal via the PDH technique Drever1983 ; Black2001 . In addition, the REFL port serves as an auxiliary monitor for the common Sagnac mode (CSAG), corresponding to the average optical path length of the clockwise (CW) and counter-clockwise (CCW) beams, and contributes to maintaining the overall optical resonance and laser-frequency stability.
IV.2.3 Pick-off port (POP)
The POP port is a pick-off of the intracavity field in the PRC, designed to sense both the common and differential components of the optical field Aston2012 . Two photodiodes are placed at positions with distinct Gouy phases to spatially separate optical modes: